# On the large-time behaviour of affine Volterra processes

### v2

@@ Abstract.
We show the existence of a stationary measure for a class of multidimensional stochastic Volterra systems of affine type. These processes are in general not Markovian, a shortcoming which hinders their large-time analysis. We circumvent this issue by lifting the system to a measure-valued stochastic PDE introduced by Cuchiero and Teichmann [15], whence we retrieve the Markov property. Leveraging on the associated generalised Feller property, we extend the Krylov-Bogoliubov theorem to this infinite-dimensional setting and thus establish an approach to the existence of invariant measures. We present concrete examples, including the rough Heston model from Mathematical Finance.
Key words and phrases:
affine Volterra, ergodic behaviour, stochastic partial differential equations, generalised Feller, rough volatility
2010 Mathematics Subject Classification:
60G22, 60F05, 60H15, 60J25
A.P. and A.J. are supported by an EPSRC EP/T032146/1 granth. K.S. is partially supported by the National Science Foundation (DMS 2107856) and Simons Foundation Award 672441. The authors would like to thank the reviewer for a very careful review of the manuscript that identified an issue with the initial version of the result and for allowing us to address it.
${}^{\ast}$ Corresponding author
@@ 1. Introduction
A different point of view is given by the theory of Random Dynamical Systems where Markovianity is replaced by a cocycle property for the driving noise, e.g. (fractional) Brownian motion [4, 33]. Furthermore, in this context one analyses the asymptotic behaviour by taking the starting time to $\infty$- ∞ and looking at the system at $t=0$, instead of starting at $t=0$ and looking at $t$ goes to +$\infty$+ ∞.
@@
The idea of [38] with augmenting the state space to recover Markovianity made its way to rough volatility modelling, albeit with a mild adaptation. Motivated by hedging applications in the rough Heston model, El Euch and Rosenbaum [25] showed that since the variance process is non-Markovian one needs to include the whole forward variance curve; this suggests to consider the system $(S_{t},(\mathbb{E}[V_{s+t}|\mathcal{F}_{t}])_{s\geq 0})$. The forward variance even satisfies a Stochastic Partial Differential Equation (SPDE) [2], which allowed to further characterise the Markovian structure of the model. A similar SPDE lift for another singular SVE was in fact derived in an earlier work [53]. The analysis of the rough Heston model is facilitated by its affine structure, unlike the rough Bergomi model, its main competitor. Yet, the forward variance curve of the latter also satisfies an SPDE as it belongs to the realm of polynomial processes [13]. The most conducive SPDE lift for our purposes, though, was introduced by Cuchiero and Teichmann in [15]. Unlike the previous approaches, they argue it is more sensible to start from the Markovian lift before solving the SVE. More precisely, they consider the one-dimensional measure-valued SPDE
@@
where $\lambda_{0}$ and $\nu$ are signed measures on $[0,+\infty]$
@@
where $\langle y,\lambda\rangle:=\int_{0}^{\infty}y(x)\,\lambda(\mathrm{d}x)$ for all $y\in\mathcal{C}_{b}(\mathbb{R}_{+})$. In fact, one should interpret the SPDE (1.2) in the mild sense such that
@@
The added value of this approach lies in the generalised Feller property, a notion introduced in [18], which is satisfied by the solution to (1.2) (albeit with $\theta=0$) and opens the gates to many parallels with Markovian ergodic theory. Indeed, it is an extension of the standard Feller property to spaces that are not locally compact. To the best of our knowledge, the present paper is the first to study the large-time behaviour of generalised Feller processes. Interestingly enough, in [15], the authors also prove weak existence and uniqueness of the mild solution of (1.2) by computing its Laplace transform
@@
As one can expect, the literature on measure-valued SPDEs is less developed. However, there exist connections with measure-valued branching processes (aka superprocesses) which are also Markovian processes characterised by their Laplace transform, only the PDE satisfied by $y$ takes a different form [51]. We note that the ergodic behaviour of superprocesses has been studied extensively thanks to their Laplace transform [11, 28, 31, 44, 47], which is a hopeful message for us. Furthermore, the analogy with our SPDE lift does not stop there: the density field of some of these superprocesses also satisfies an SPDE with an affine structure [48, 52], and in the continuous-state case branching processes share properties with affine SDEs [45]. However, superprocesses take values in spaces of non-negative measures, which can be locally compact, in which case it makes sense to use the standard Feller property.
The rest of the paper is organised as follows. In Section 2 we discuss the mathematical framework, the generalised Feller property, and present our main results, Proposition 2.4 and Theorem 2.7. In Section 3 we introduce a condition for existence of an invariant measure in an abstract setting that then leads to the proof of Proposition 2.4. In Section 4, we consider the multidimensional SPDE, prove that there is a unique solution to the multidimensional lift that is a generalised Feller process and further discuss its properties in Proposition 4.3. Section 5 concludes the proof of Theorem 2.7 by showing that under Assumption 2.6, the bound in Proposition 2.4 holds. Section 6 discusses future directions.
@@ 2. Framework and main results
We recall some of the notations introduced in [15]. Let $\bm{X}$ be a completely regular Hausdorff topological space. We say that $\varrho\colon\bm{X}\to(0,\infty)$ is an admissible weight function if the sets $K_{R}:=\{x\in\bm{X}:\varrho(x)\leq R\}$ are compact for all $R>0$ [15, Definition 2.1]. The supremum norm on $\mathbb{R}$ is denoted $\left\lVert\cdot\right\rVert_{\infty}$. The vector space
@@
is a Banach space, and $\mathcal{C}_{b}(\bm{X})\subset\mathrm{B}^{\varrho}(\bm{X})$. Moreover, the space $\mathcal{B}^{\varrho}(\bm{X})$ is defined as the closure of $\mathcal{C}_{b}(\bm{X})$ in $\mathrm{B}^{\varrho}(\bm{X})$ [15, Definition 2.3], and is also a Banach space when equipped with the norm (2.1). Generalised Feller semigroups are the analogue of standard Feller semigroups on the space of continuous functions vanishing at infinity on locally compact spaces. They are bounded, positive, linear, strongly continuous operators.
@@
Theorems 2.11 and 2.13 of [15] ensure that each generalised Feller semigroup gives rise to an associated Markov process and such process has a version with càglàd paths. Let $\mathcal{P}(\bm{X})$ be the space of probability measures on $\bm{X}$.
@@ 2.1. Main results
$\sup_{t\geq 0}\mathbb{E}[\left\lVert\lambda_{t}\right\rVert_{Y^{\star}}]<\infty,\vspace{-5pt}$ (2.2)
@@
Our goal is to apply this to $Y^{\star}$-valued SPDEs which we introduce now. Let $d\geq 1$, we consider the Banach space $Y:=\mathcal{C}_{b}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ (compactifying $\mathbb{R}_{+}$ makes $\mathcal{C}_{b}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ separable) and its dual $Y^{\star}=\mathcal{M}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ is the space of signed measured, as explained in [20, Chapter IV]. We then have $\langle y,\lambda\rangle:=\sum_{i=1}^{d}\langle y^{i},\lambda^{i}\rangle_{i}:=\sum_{i=1}^{d}\int_{0}^{\infty}y^{i}(x)\lambda^{i}(\mathrm{d}x)$ for all $y\in Y$, $\lambda\in Y^{\star}$. Moreover, the weight function is $\varrho(\lambda):=1+\left\lVert\lambda\right\rVert_{Y^{\star}}$, the set $\bm{X}$ introduced in [15, Section 1.3] is a subset of $Y^{\star}$ and the space $\mathcal{B}^{\varrho}(\bm{X})$ is defined in the same way.
We consider the $Y^{\star}$-valued multidimensional SPDE
@@
    $\beta_{ii}>0$ for $i=1,\cdots,\tilde{d}$ and $\beta_{ik}\geq 0$ for $j,k=1,\cdots,\tilde{d}$ and $j\neq k$.
@@
Assumption 2.6 has specific structural restrictions on the coefficients $c,\beta,\sigma$ under which Theorem 2.7 holds. We note here though that these are only sufficient conditions and in fact, intermediate results of independent interest, such as Proposition 4.3 and Lemma 5.2, do not require these conditions.
@@
$\mathrm{d}X^{i}_{t}$ $=-\beta_{ii}\overline{\lambda}^{i}_{t}\,\mathrm{d}t+\sqrt{\overline{\lambda}^{i}_{t}}\sum_{j=1}^{m}\sigma_{iji}\mathrm{d}W^{j}_{t},\qquad\qquad\qquad$ if $\text{if }i\in\llbracket 1,\tilde{d}\rrbracket,$
$\mathrm{d}X^{i}_{t}$ $=-\sum_{k=1}^{\tilde{d}}\beta_{ik}\overline{\lambda}^{k}_{t}\,\mathrm{d}t+\sum_{j=1}^{m}\sum_{k=1}^{\tilde{d}}\sigma_{ijk}\sqrt{\overline{\lambda}_{t}^{k}}\,\mathrm{d}W^{j}_{t},$ if $\text{if }i\in\llbracket\tilde{d}+1,d\rrbracket.$
@@
    Its lift is the one-dimensional SPDE introduced in [15].
@@
Under Assumption 2.6, there exists a probability measure $\mu^{\star}$ on $Y^{\star}$ such that if, for any $s\geq 0$, $\mu^{\star}$ is the distribution of $\lambda_{s}$ then $(\overline{\lambda}_{t})_{t\geq s}$ is a strictly stationary process on $\mathbb{R}_{+}$.
@@
  • In the closest paper to ours [32], the authors prove that $(\overline{\lambda}_{t+h})_{t\geq 0}$ converges towards a stationary process as $h\to\infty$. They characterise properties of the marginal law of this limiting process but do not provide information on its dynamics. Moreover, this convergence result does not imply that $\overline{\lambda}$ itself is stationary.
@@
Based on integral representations for the Mittag-Leffler function [43, Section 7], we obtain
@@ 3. A condition for existence
This section aims at proving Proposition 2.4. We start by stating and proving an extension of Krylov-Bogoliubov theorem [16, Theorem 11.7] to the setting of generalised Feller processes on $\mathcal{B}^{\varrho}(\bm{X})$ for any completely regular Hausdorff topological space $\bm{X}$.
@@
Let $(P_{t})_{t\geq 0}$ be a generalised Feller semigroup. Suppose that there exists $\gamma\in\mathcal{P}(\bm{X})$ and a strictly positive sequence $T_{n}$ going to +$\infty$+ ∞ as $n$ goes to +$\infty$+ ∞ such that
@@
    for any $\bm{X}$-valued random variable $\lambda$ with distribution $\gamma$,
    $\sup_{t\geq 0}\int_{\bm{X}}P_{t}\varrho(\lambda)\gamma(\mathrm{d}\lambda),<\infty$ (3.1)
@@
By Prokhorov’s theorem [16, Theorem 2.3], (i) is equivalent to tightness of $(P_{t}^{\ast}\gamma)_{t\geq 0}$. The original Krylov-Bogoliubov theorem for standard Feller semigroups does not require (ii).
@@
Fix $t>0$ and $\varphi\in\mathcal{B}^{\varrho}(\bm{X})$, then by definition $P_{t}\varphi\in\mathcal{B}^{\varrho}(\bm{X})$. Let $R>0$, since $f\in\mathcal{B}^{\varrho}(\bm{X})$ if and only if $f\lvert_{K_{R}}\in\mathcal{C}(K_{R})$ (see [15, Equation (2.3)]), we need to work on $K_{R}$ to apply the weak convergence of (i):
@@
Then we go back to $\bm{X}$ to apply the adjoint property
@@
Note that in the last computation, weak convergence of $Q_{T_{n}}^{\ast}\gamma$ to $\mu$ can be used even though the function $\mathbbm{1}_{K_{R}}\varphi$ is not in $\mathcal{C}(\bm{X})$. This can be done following the proof of approximation by appropriate cut-off functions as in [6, Theorem 29.1]. Overall, this yields our objective plus a remainder
@@
To deal with $\varepsilon(R)$, we apply a Fatou-type lemma for measures [21, Theorem A.3.12]. For any $f\in\mathcal{B}^{\varrho}(\bm{X})$,
@@
The same computations also yield the same estimate for $\left\lvert\int_{\bm{X}\setminus K_{R}}f\,\mathrm{d}Q_{T_{n}}^{\ast}\gamma\right\rvert$. By [15, Equation (2.3)] we have for any $f\in\mathcal{B}^{\varrho}(\bm{X})$,
@@
The uniform bound (3.1) thus ensures that $\varepsilon(R)$ vanishes as $R$ goes to +$\infty$+ ∞. ∎
@@
Furthermore, by [8, Lemma 2.9], for all $t\geq 0$, let $\Lambda_{t}$ be the $Y^{\star}$-valued random variable with distribution $Q^{\ast}_{t}\gamma$. The family $(\Lambda_{t})_{t\geq 0}$ is tight if and only if there exists a tightness function $G:Y^{\star}\to\mathbb{R}_{+}$ such that $\sup_{t\geq 0}\mathbb{E}[G(\Lambda_{t})]<\infty$. We observe that an admissible weight function is also a tightness function by design, hence proving tightness of $(\Lambda_{t})_{t\geq 0}$ boils down to showing
@@
which is a consequence of (3.3). Therefore, both conditions of Lemma 3.1 are satisfied as soon as (2.2) holds. ∎
4. The multidimensional SPDE

### v3

@@ Abstract.
We show the existence of a stationary measure for a class of multidimensional stochastic Volterra systems of affine type. These processes are in general not Markovian, a shortcoming which hinders their large-time analysis. We circumvent this issue by lifting the system to a measure-valued stochastic evolution equation introduced by Cuchiero and Teichmann [17], whence we retrieve the Markov property. Leveraging on the associated generalised Feller property, we extend the Krylov-Bogoliubov theorem to this infinite-dimensional setting and thus establish an approach to the existence of invariant measures. We present concrete examples, including the rough Heston model from Mathematical Finance.
@@ 1. Introduction
A different point of view is given by the theory of Random Dynamical Systems where Markovianity is replaced by a cocycle property for the driving noise, e.g. (fractional) Brownian motion [4, 35]. Furthermore, in this context one analyses the asymptotic behaviour by taking the starting time to -$\infty$ and looking at the system at $t=0$, instead of starting at $t=0$ and looking at $t$ goes to $\infty$.
@@
The idea of [40] with augmenting the state space to recover Markovianity made its way to rough volatility modelling, albeit with a mild adaptation. Motivated by hedging applications in the rough Heston model, El Euch and Rosenbaum [27] showed that since the variance process is non-Markovian one needs to include the whole forward variance curve; this suggests to consider the system $(S_{t},(\mathbb{E}[V_{s+t}|\mathcal{F}_{t}])_{s\geq 0})$. The forward variance even satisfies a Stochastic Partial Differential Equation (SPDE) [2], which allowed to further characterise the Markovian structure of the model. A similar stochastic evolution equation lift for another singular SVE was in fact derived in an earlier work [55]. The analysis of the rough Heston model is facilitated by its affine structure, unlike the rough Bergomi model, its main competitor. Yet, the forward variance curve of the latter also satisfies a stochastic evolution equation as it belongs to the realm of polynomial processes [15]. The most conducive stochastic evolution equation lift for our purposes, though, was introduced by Cuchiero and Teichmann in [17]. Unlike the previous approaches, they argue it is more sensible to start from the Markovian lift before solving the SVE. More precisely, they consider the one-dimensional measure-valued stochastic evolution equation
@@
where $\lambda_{0}$ and $\nu$ are signed measures on $0,+\infty]$
@@
where $\langle y,\lambda\rangle:=\int_{0}^{\infty}y(x)\,\lambda(\mathrm{d}x)$ for all $y\in\mathcal{C}_{b}(\mathbb{R}_{+})$. In fact, one should interpret the stochastic evolution equation (1.2) in the mild sense such that
@@
The added value of this approach lies in the generalised Feller property, a notion introduced in [20], which is satisfied by the solution to (1.2) (albeit with $\theta=0$) and opens the gates to many parallels with Markovian ergodic theory. Indeed, it is an extension of the standard Feller property to spaces that are not locally compact. To the best of our knowledge, the present paper is the first to study the large-time behaviour of affine Volterra processes through generalised Feller processes. Friesen and Karbach, in a paper posted at the same time as ours [34], exploit generalised Feller processes to study the stationarity of a class of Hilbert-valued models. Interestingly enough, in [17], the authors also prove weak existence and uniqueness of the mild solution of (1.2) by computing its Laplace transform
@@
As one can expect, the literature on measure-valued stochastic evolution equations is less developed. However, there exist connections with measure-valued branching processes (aka superprocesses) which are also Markovian processes characterised by their Laplace transform, only the PDE satisfied by $y$ takes a different form [53]. We note that the ergodic behaviour of superprocesses has been studied extensively thanks to their Laplace transform [13, 29, 32, 46, 49], which is a hopeful message for us. Furthermore, the analogy with our stochastic evolution equation lift does not stop there: the density field of some of these superprocesses also satisfies an stochastic evolution equation with an affine structure [50, 54], and in the continuous-state case branching processes share properties with affine SDEs [47]. However, superprocesses take values in spaces of non-negative measures, which can be locally compact, in which case it makes sense to use the standard Feller property.
The rest of the paper is organised as follows. In Section 2 we discuss the mathematical framework, the generalised Feller property, and present our main results, Proposition 2.4 and Theorem 2.7. In Section 3 we introduce a condition for existence of an invariant measure in an abstract setting that then leads to the proof of Proposition 2.4. In Section 4, we consider the multidimensional stochastic evolution equation, prove that there is a unique solution to the multidimensional lift that is a generalised Feller process and further discuss its properties in Proposition 4.3. Section 5 concludes the proof of Theorem 2.7 by showing that under Assumption 2.6, the bound in Proposition 2.4 holds. Section 6 discusses future directions.
@@ 2. Framework and main results
We recall some of the notations introduced in [17]. Let $\bm{X}$ be a completely regular Hausdorff topological space. We say that $\varrho\colon\bm{X}\to(0,\infty)$ is an admissible weight function if the sets $K_{R}:=\{x\in\bm{X}:\varrho(x)\leq R\}$ are compact for all $R>0$ [17, Definition 2.1]. The supremum norm on $\mathbb{R}$ is denoted $\left\lVert\cdot\right\rVert_{\infty}$. The vector space
@@
is a Banach space, and $\mathcal{C}_{b}(\bm{X})\subset\mathrm{B}^{\varrho}(\bm{X})$. Moreover, the space $\mathcal{B}^{\varrho}(\bm{X})$ is defined as the closure of $\mathcal{C}_{b}(\bm{X})$ in $\mathrm{B}^{\varrho}(\bm{X})$ [17, Definition 2.3], and is also a Banach space when equipped with the norm (2.1). Generalised Feller semigroups are the analogue of standard Feller semigroups on the space of continuous functions vanishing at infinity on locally compact spaces. They are bounded, positive, linear, strongly continuous operators.
@@
Theorems 2.11 and 2.13 of [17] ensure that each generalised Feller semigroup gives rise to an associated Markov process and such process has a version with càglàd paths.
Let $\mathcal{P}(\bm{X})$ be the space of probability measures on $\bm{X}$.
@@ 2.1. Main results
$\sup_{t\geq 0}\mathbb{E}[\left\lVert\lambda_{t}\right\rVert_{Y^{\star}}]<\infty,\vskip-5.0pt$ (2.2)
@@
Our goal is to apply this to $Y^{\star}$-valued stochastic evolution equations which we introduce now. Let $d\geq 1$, we consider the Banach space $Y:=\mathcal{C}_{b}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$, where $\overline{\mathbb{R}}_{+}:=\mathbb{R}\cup{+\infty}$ (compactifying $\mathbb{R}_{+}$ makes $\mathcal{C}_{b}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ separable) and its dual $Y^{\star}=\mathcal{M}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ is the space of signed measured, as explained in [22, Chapter IV]. We then have $\langle y,\lambda\rangle:=\sum_{i=1}^{d}\langle y^{i},\lambda^{i}\rangle_{i}:=\sum_{i=1}^{d}\int_{0}^{\infty}y^{i}(x)\lambda^{i}(\mathrm{d}x)$ for all $y\in Y$, $\lambda\in Y^{\star}$. Moreover, the weight function is $\varrho(\lambda):=1+\left\lVert\lambda\right\rVert_{Y^{\star}}$, the set $\bm{X}$ introduced in [17, Section 1.3] is a subset of $Y^{\star}$ and the space $\mathcal{B}^{\varrho}(\bm{X})$ is defined in the same way.
We consider the $Y^{\star}$-valued multidimensional measure-valued stochastic evolution equation
@@
    $\beta_{ii}>0$ for $i=1,\cdots,\tilde{d}$ and $\beta_{jk}\leq 0$ for $j,k=1,\cdots,\tilde{d}$ and $j\neq k$.
@@
Assumption 2.6 has specific structural restrictions on the coefficients $\beta,\sigma$ under which Theorem 2.7 holds. We note here though that these are only sufficient conditions and in fact, intermediate results of independent interest, such as Proposition 4.3 and Lemma 5.2, do not require these conditions. We also note that even though our main result holds with $c_{ijk}=0$ for all $i,j,k$ in (2.5), certain results in the paper are proven with non-zero $c_{ijk}$, see Section 4.
@@
$\mathrm{d}X^{i}_{t}$ $=-\beta_{ii}\overline{\lambda}^{i}_{t}\,\mathrm{d}t+\sqrt{\overline{\lambda}^{i}_{t}}\sum_{j=1}^{m}\sigma_{iji}\mathrm{d}W^{j}_{t},\qquad\qquad\qquad$ $\text{if }i\in\llbracket 1,\tilde{d}\rrbracket$,
$\mathrm{d}X^{i}_{t}$ $=-\sum_{k=1}^{\tilde{d}}\beta_{ik}\overline{\lambda}^{k}_{t}\,\mathrm{d}t+\sum_{j=1}^{m}\sum_{k=1}^{\tilde{d}}\sigma_{ijk}\sqrt{\overline{\lambda}_{t}^{k}}\,\mathrm{d}W^{j}_{t},$ $\text{if }i\in\llbracket\tilde{d}+1,d\rrbracket.$
@@
    Its lift is the one-dimensional stochastic evolution equation introduced in [17].
@@
Under Assumption 2.6, there exists a probability measure $\mu^{\star}$ on $Y^{\star}=\mathcal{M}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$ such that if, for any $s\geq 0$, $\mu^{\star}$ is the distribution of $\lambda_{s}$ then $(\overline{\lambda}_{t})_{t\geq s}$ is a strictly stationary process on $\mathbb{R}_{+}$.
@@
  • In the closest paper to ours [33], the authors prove that $(\overline{\lambda}_{t+h})_{t\geq 0}$ converges in law towards a stationary process $(\overline{\lambda}_{t}^{{\rm stat}})_{t\geq 0}$ as $h\to\infty$. They characterise properties of the marginal law of this limiting process but do not provide information on its dynamics, in particular it is not known whether $\overline{\lambda}^{{\rm stat}}$ satisfies an SVE. Moreover, this convergence result does not imply that $\overline{\lambda}$ itself is a stationary process.
@@
Based on integral representations for the Mittag-Leffler function [45, Section 7], we obtain
@@ 3. A condition for existence
This section aims at proving Proposition 2.4. We start by stating and proving an extension of Krylov-Bogoliubov theorem [18, Theorem 11.7] to the setting of generalised Feller processes on $\mathcal{B}^{\varrho}(\bm{X})$ for any completely regular Hausdorff topological space $\bm{X}$.
@@
Let $(P_{t})_{t\geq 0}$ be a generalised Feller semigroup. Suppose that there exists $\gamma\in\mathcal{P}(\bm{X})$ and a strictly positive sequence $T_{n}$ going to $\infty$ as $n$ goes to $\infty$ such that
@@
    It holds that
    $\sup_{t\geq 0}\int_{\bm{X}}P_{t}\varrho(\lambda)\gamma(\mathrm{d}\lambda)<\infty$ (3.1)
@@
The original Krylov-Bogoliubov theorem for standard Feller semigroups does not require (ii).
On Polish spaces, Prokhorov’s theorem [18, Theorem 2.3] states that (i) is equivalent to tightness of $(P_{t}^{\ast}\gamma)_{t\geq 0}$. We will invoke an extension of this result to completely regular spaces [8, Chapitre IX, Section 5.5, Théorème 1] as dual spaces equipped with the weak-$\star$-topology are not Polish. This result relies upon the so-called Prokhorov condition for a subset $H$ of the space of bounded Radon measures $\mathcal{M}_{b}(T)$, where $T$ is completely regular. In the case $H=\mathcal{P}(\bm{X})\subset\mathcal{M}_{b}(\bm{X})$ of interest to us, this preliminary condition is equivalent to tightness. We now state the theorem.
Theorem 3.3 (Bourbaki).
Let $T$ be a completely regular space, and $H$ a subset of $\mathcal{M}_{b}(T)$ satisfying Prokhorov’s condition; then $H$ is relatively compact in $\mathcal{M}_{b}(T)$ equipped with the topology of weak convergence.
As the dual of a separable Banach space equipped with the weak-$\ast$-topology, $Y^{\star}$ is a completely regular space and so is the subset $\bm{X}$.
@@
Fix $t>0$ and $\varphi\in\mathcal{B}^{\varrho}(\bm{X})$, then by definition $P_{t}\varphi\in\mathcal{B}^{\varrho}(\bm{X})$. Let $R>0$, we know from [17, Equation (2.3)] that $f\in\mathcal{B}^{\varrho}(\bm{X})$ if and only if $f\lvert_{K_{R}}\in\mathcal{C}(K_{R})$, where we recall $K_{R}=\{x\in\bm{X}:\varrho(x)\leq R\}$. We thus need to work on $K_{R}$ to apply the weak convergence of (i):
@@
Note that in the last computation, weak convergence of $Q_{T_{n}}^{\ast}\gamma$ to $\mu$ can be used even though the function $\mathbbm{1}_{K_{R}}\varphi$ is not in $\mathcal{C}(\bm{X})$. This can be done following the proof of approximation by appropriate cut-off functions as in [7, Theorem 29.1]. Then we go back to $\bm{X}$ to apply the adjoint property
@@
In the last computation, once again the weak convergence of $Q_{T_{n}}^{\ast}\gamma$ to $\mu$ can be used even though the function $\mathbbm{1}_{K_{R}}\varphi$ is not in $\mathcal{C}(\bm{X})$, as in [7, Theorem 29.1]. Overall, this yields our objective plus a remainder
@@
To deal with $\varepsilon(R)$, we apply a Fatou-type lemma for measures [23, Theorem A.3.12]. For any $f\in\mathcal{B}^{\varrho}(\bm{X})$,
@@
The same computations also yield the same estimate for $\left\lvert\int_{\bm{X}\setminus K_{R}}f\,\mathrm{d}Q_{T_{n}}^{\ast}\gamma\right\rvert$. By [17, Equation (2.3)] we have for any $f\in\mathcal{B}^{\varrho}(\bm{X})$,
@@
The uniform bound (3.1) thus ensures that $\varepsilon(R)$ vanishes as $R$ goes to $\infty$. ∎
@@
For all $t\geq 0$, let $\Lambda_{t}$ be the $Y^{\star}$-valued random variable with distribution $Q^{\ast}_{t}\gamma$. Inspection of the proof of [10, Lemma 2.9] shows that, even though $Y^{\star}$ is not a Polish space, the family $(\Lambda_{t})_{t\geq 0}$ is tight if there exists a tightness function $G:Y^{\star}\to\mathbb{R}_{+}$ such that $\sup_{t\geq 0}\mathbb{E}[G(\Lambda_{t})]<\infty$. We observe that an admissible weight function is also a tightness function by design, hence proving tightness of $(\Lambda_{t})_{t\geq 0}$ boils down to showing
@@
which is a consequence of (3.3). We then invoke the version of Prokhorov’s theorem on completely regular spaces (Theorem 3.3) to deduce that $(Q^{\ast}_{t}\gamma)_{t\geq 0}$ is relatively compact in $\mathcal{P}(\bm{X})$ and hence has a converging subsequence.
Therefore, both conditions of Lemma 3.1 are satisfied as soon as (2.2) holds. ∎
4. The multidimensional measure-valued stochastic evolution equation